OpenStudy (anonymous):

Write an equation to represent the pattern in the table below. h = 10s + 50 s = 10h +50 s = 10h – 50 h = 10h -50

OpenStudy (anonymous):

The graph

OpenStudy (anonymous):

h=10s+50 5=10(100)+50 5=1000+50 5=1050 s=10h+50 100=10*5+50 100=50+50 100=100 s=10h-50 100=10*5-50 100=50-50 100=0 h=10h-50 5=10*100-50 5=1000-50 5=950

OpenStudy (anonymous):

So it's s= 10h +50

OpenStudy (anonymous):

yes